# UnZipLoRA: Separating Content and Style from a Single Image

### v1

@@ Abstract
* Currently at Google.   Currently at Meta.
@@
This paper introduces UnZipLoRA, a method for decomposing an image into its constituent subject and style, represented as two distinct LoRAs (Low-Rank Adaptations). Unlike existing personalization techniques that focus on either subject or style in isolation, or require separate training sets for each, UnZipLoRA disentangles these elements from a single image by training both the LoRAs simultaneously. UnZipLoRA ensures that the resulting LoRAs are compatible, i.e., they can be seamlessly combined using direct addition. UnZipLoRA enables independent manipulation and recontextualization of subject and style, including generating variations of each, applying the extracted style to new subjects, and recombining them to reconstruct the original image or create novel variations. To address the challenge of subject and style entanglement, UnZipLoRA employs a novel prompt separation technique, as well as column and block separation strategies to accurately preserve the characteristics of subject and style, and ensure compatibility between the learned LoRAs. Evaluation with human studies and quantitative metrics demonstrates UnZipLoRA’s effectiveness compared to other state-of-the-art methods, including DreamBooth-LoRA, Inspiration Tree, and B-LoRA.
@@ 1 Introduction
Imagine an artist inspired by a single image – perhaps a microscope rendered in a unique illustration technique (Fig. 1). This image depicts both the content or subject (the microscope with its specific shape and details) and style (the characteristic artistic technique). What if we could disentangle these intertwined elements, enabling the artist to extract and manipulate them independently? This capability, long sought in the computer vision literature [36], would open up new avenues for artistic expression, allowing for the destylization and recontextualization of the subject in different conditions, the application of the extracted style to new subjects, and the creation of entirely novel combinations while preserving both the original subject and style.
@@
The recent surge of diffusion models [13, 35, 27] has unlocked previously unprecedented automatic image generation capabilities. The primary means of controlling such models is through text prompts, but text-based conditioning is inadequate for capturing the details of nuanced concepts, such as specific object instances (my dog) or individual styles (my child’s crayon drawing). Example-driven generation is highly desired in such scenarios. To this end, model fine-tuning methods like DreamBooth [29] and StyleDrop [33] captures subject or style of the reference image(s), and allow for the generation of novel renditions of specific subjects or styles. However, such approaches tend to focus on either content or style in isolation, and cannot be easily made to capture both or perform disentanglement. On the other hand, stylization techniques like ZipLoRA [32] and B-LoRA [6] can combine the subject and style together, but require the training of two separate models using subject- and style-specific input images. However, scenario we consider requires the opposite,i.e., decomposing the subject and style from a single input image.
@@
This challenge of separating content and style from a single image is precisely what we tackle in this work. We introduce UnZipLoRA, a novel method that deconstructs an image into its constituent subject and style, represented as two distinct LoRAs (Low-Rank Adaptations [14]) trained simultaneously. These LoRAs can be used independently to generate variations of the subject or style, and allow for recontextualizations. Moreover, our joint training method ensures that the resulting LoRAs are inherently compatible,i.e., can be seamlessly combined by direct addition to reconstruct the original image or to generate novel compositions of subject and style while preserving their fidelity. In fact our approach can be seen as a next stage of concept extraction - a problem previously studied in concept decomposition methods like Inspiration Tree [37] and CusConcept [40] that rely on textual inversion [8] to learn multiple text embeddings corresponding to the subconcepts within a set of images. However, textual inversion alone, without weight fine-tuning, does not provide adequate control over or fidelity of the extracted concepts that remain generic in nature and do not capture nuances of the input object/style.
@@
Our key challenge is to learn two independent LoRAs simultaneously using only a single input image as supervision while ensuring that the resulting LoRAs correctly capture the subject and style concepts. Typical LoRA fine-tuning operates by binding the LoRA weights with the trigger phrase representing a specific subject or style concept in the input image – such as “a <$c$> dog in <$s$> painting style”. If we apply such a naive approach in our case, the presence of both the trigger phrases in a single input prompt makes it difficult for the subject and style LoRAs to bind to the correct trigger phrase, resulting in cross-contamination. To solve this problem, we propose a novel prompt separation strategy that uses different prompts for each LoRA and the base model, and then combines them together in the intermediate feature space of the diffusion model in such a way that the loss for each LoRA can be calculated jointly using only the input image as supervision.
While prompt separation allows for joint training of LoRAs, the resulting LoRAs may not be compatible with each other, i.e., combining them through direct addition may produce poor quality recontextualizations. To make them inherently compatible, we also propose column separation and block separation strategies built upon insights from ZipLoRA [32] and B-LoRA [6] respectively. Column separation determines the importance of each column of LoRA weight matrix and dynamically assigns each column to either a subject LoRA or style LoRA. Such disjoint assignment ensures that high-importance columns from each LoRA are not added together. Block separation reserves some blocks of the U-net predominantly for style or for the subject, ensuring a further degree of disentanglement.
We demonstrate the effectiveness of UnZipLoRA for accurate separation of content and style using both human studies and quantitative metrics. Our results show a clear advantage of our method over separate LoRA fine-tuning via DreamBooth [29], concept separation via Inspiration Tree [37], and even the most recent state-of-the-art B-LoRA method [6]. We also showcase our method’s ability to preserve the concept fidelity for wide array of recontextualizations – whether for using subject or style separately or together.
@@ 2 Related Work
Fine-tuning diffusion models is an effective way to personalize the text-to-image (T2I) models to depict specific concepts based on textual descriptions. Textual Inversion [8] optimizes the text embedding to represent a specific visual concept. DreamBooth [29] fine-tunes the diffusion model itself to better capture an input concept from a small number of images. Similarly, another group of methods [10, 17] aims to optimize specific parts of the networks to capture visual concepts. Importantly, LoRA [14] is a finetuning techique that only optimizes a small subset of weights by low-rank approximations. Personalization approaches quickly adopt this techique, as it is efficient for training and can mitigate overfitting problems.
@@
Image stylization is an area of research dating back at least 20 years [5, 12]. Great advances in arbitrary style transfer was achieved by the convolutional neural network-based approaches [9, 16, 15, 19, 24]. With the advent of image generative models, a variety of approaches have attempted to fine-tune a pre-trained generative model such as GAN or diffusion model for the task of stylization [20, 7, 43, 23, 38, 18, 3, 31, 42, 41, 33, 4]. While these works provide valuable insights into style learning using generative models, the task we attempt to solve in this work is the opposite: instead of stylizing a content image, we attempt to decompose a stylistic image into its subject and style counterparts.
@@
Content-style decomposition. The task of subject-style decomposition can also be seen as a type of concept extraction. Inspiration Tree [37] aims to learn multiple embeddings corresponding to hierarchical subconcepts within a set of images. However, its reliance on textual inversion [8] limits this method to primarily producing text embeddings rather than the full LoRA weights needed for granular control over generation. Lego [22] is another method based on textual inversion that explores inverting personalized concepts from images that go beyond object appearance and style. Similarly, CusConcept [40] decomposes an image into visual concepts by learning customized embeddings. However, it does not produce dedicated LoRA models for each concept. This limitation, along with its reliance on computationally expensive LLMs and its lack of explicit content-style decomposition, restricts its practical applicability and flexibility. U-VAP [39] is a fine-tuning method that allows users to specify desired attributes from a set of images, enabling the disentangled use of visual concepts in diverse settings. However, this method requires elaborate data augmentation aided by LLMs, posing scalability challenges. ConceptExpress [11] explores unsupervised concept extraction and recreation by leveraging the inherent capabilities of pre-trained diffusion models. However, its reliance on localized masks extracted from the U-net limits its application to concepts that can be localized within an image. This constraint prevents it from effectively handling global concepts like style. This limitation applies to Break-A-Scene [1] too, which relies on localized masks.
@@
The method most directly aimed at content and style separation is B-LoRA [6]. The authors of this work analyze the architecture of the SDXL base model [25] to find blocks most responsible for capturing content and style. By jointly training content and style LoRAs restricted to the respective blocks, they obtain models that can be successfully mixed and matched for various stylization applications. We take advantage of the same block constraints in our work, but find that they are not sufficient. Tellingly, B-LoRA [6] does not demonstrate results for de-stylization – i.e., given a stylized image, render the same subject in a realistic style. In our experiments, we find that B-LoRA struggles to extract and represent subjects in their unstylized appearance, hindering its ability to faithfully manipulate content. The prompt separation strategy of UnZipLoRA provides a significant improvement in disentanglement ability, and column separation further enhances the fidelity of fine details.
@@ 3 Method
Figure 2: Overview. (1): Our training pipeline takes three prompts to learn two separate LoRAs. The column masks are introduced to establish the orthogonality. (2)-(4): Inference with individual learned LoRA or the combined LoRA.
@@ 3.1 Preliminaries
Diffusion models.  [13, 35, 27] represent the current state-of-the-art in generative modeling, renowned for their ability to synthesize high-quality, photorealistic images. The forward diffusion process gradually adds Gaussian noise to an image until it becomes indistinguishable from random noise, while the training stage learns a model for the reverse process that reconstructs the original image by gradually removing noise. This reconstruction is typically achieved using a U-net architecture with text conditioning, enabling text-to-image generation. In this work, we focus on latent diffusion models (LDMs) [27], which operate in a lower-dimensional latent space. Specifically, we utilize Stable Diffusion XL v1 [25] for all our experiments.
LoRA fine-tuning. LoRA (Low-Rank Adaptation) is a method for efficient adaptation of large vision or language models to a new downstream task [14, 30]. The key concept of LoRA is that the weight updates $\Delta W$ to the base model weights $W_{0}\in\mathbb{R}^{m\times n}$ during fine-tuning have a low intrinsic rank, thus the update $\Delta W$ can be decomposed into two low-rank matrices $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ for efficient parameterization with $\Delta W=BA$. Here, $r$ represents the intrinsic rank of $\Delta W$ with $r<<min(m,n)$. During training, only $A$ and $B$ are updated to find suitable $\Delta W=BA$, while keeping $W_{0}$ constant. For inference, the updated weight matrix $W$ can be obtained as $W=W_{0}+BA$. Due to its efficiency, LoRA is widely used for fine-tuning open-sourced diffusion models. In case of text-to-image LDM models, model customization can be achieved by fine-tuning the LoRA to minimize reconstruction loss $\mathcal{L}_{DB}$ as proposed by DreamBooth [29].
@@ 3.2 Problem Setup
We aim to decouple content and style from a single input image by learning two distinct LoRAs simultaneously. Given a pre-trained diffusion model with weights $\{W_{0}^{i}\}$, we learn two models: content LoRA $L_{c}=\{\Delta W_{c}^{i}\}$ and style LoRA $L_{s}=\{\Delta W_{s}^{i}\}$. Here, $i$ denotes the index of layers of the diffusion U-net. We use the trigger phrases <$c$> and <$s$> in our text prompts to describe the content and style respectively. Once trained, the resulting LoRAs can be used either separately or together to achieve various recontextualizations as depicted in Fig. 2. For simplicity, we drop the superscript $i$ since our method operates over all the LoRA-enabled weight matrices of the base model.
@@ 3.3 UnZipLoRA
In this section, we propose our UnZipLoRA approach, which relies on three key components to ensure accurate disentanglement. First, prompt separation is introduced in Sec. 3.3.1 to train the two LoRAs with different prompts emphasizing style or content. Second, column separation is introduced in Sec. 3.3.2 to enforce the orthogonality of weights in the two LoRAs with the help of a sparse column map. Third, we introduce block separation in Sec. 3.3.3 to assign more weight to the style LoRA and content LoRA in the style-sensitive or content-sensitive blocks, respectively.
@@ 3.3.1 Prompt Separation
Our core challenge lies in the utilization of a single image as supervision for training two separate LoRAs. Since the input image contains both subject and style, we must use both LoRAs $L_{c}$ and $L_{s}$ together to calculate the loss during training. If we also use <$c$> and <$s$> in the input prompt, as in “A <$c$> in <$s$> style”, the training would lead to cross-contamination since the tokens corresponding to the style descriptor <$s$> would be attended to by the cross-attention layers of the subject LoRA, and vice versa. This is evident in Fig. 3, where both the subject and style LoRAs are overfitted to input image when a naive DreamBooth-LoRA is used.
@@
If we add the content and style LoRAs to the base model and use both <$c$> and <$s$> in the prompt $x$, the mapping in the cross-attention layer becomes
@@
where $x$ is the embedding for the combined prompt “A <$c$> in <$s$> style”, while $x_{c}$ and $x_{s}$ are the embeddings of subject and style descriptors <$c$> and <$s$> respectively. This allows each LoRA to attend to different concepts, preventing cross-contamination as shown in Fig. 3: prompt separation successfully destylizes the content along with bringing improvements in style learning.
Figure 3: Effectiveness of each component of our approach. We use DreamBooth to train two LoRAs with Eq. 2 as our baseline. The prompt separation achieves successful decomposition of the concepts into different LoRAs. The column separation disentangles the two LoRAs to make them compatible for joint inference. The block separation further boosts the quality of the learned style.

### v2

@@ Abstract
* Currently at Google.   Currently at Amazon.
@@
This paper introduces UnZipLoRA, a method for decomposing an image into its constituent subject and style, represented as two distinct LoRAs (Low-Rank Adaptations). Unlike existing techniques that focus on either subject or style in isolation, or require separate training sets for each, UnZipLoRA disentangles these elements from a single image by training both the LoRAs simultaneously. UnZipLoRA ensures that the resulting LoRAs are compatible, i.e., can be seamlessly combined using direct addition. UnZipLoRA enables independent manipulation and recontextualization of subject and style – generating variations of each, applying the extracted style to new subjects, and recombining them to reconstruct the original image or create novel variations. To address the challenge of subject and style entanglement, UnZipLoRA employs a novel prompt separation technique, as well as column and block separation strategies to accurately preserve the characteristics of subject and style and ensure compatibility between the learned LoRAs. Evaluation with human studies and quantitative metrics demonstrates UnZipLoRA’s effectiveness compared to other state-of-the-art methods, including DreamBooth-LoRA, Inspiration Tree, and B-LoRA.
@@ 1 Introduction
Imagine an artist inspired by a single image – perhaps a microscope rendered in a unique illustration technique (Fig. 1). The image depicts both the content or subject (the microscope with its specific shape and details) and style (the characteristic artistic technique). What if we could disentangle these intertwined elements, enabling the artist to extract and manipulate them independently? This capability, long sought in the computer vision literature [37], would open up new avenues for artistic expression, allowing for the destylization and recontextualization of the subject in different conditions, the application of the extracted style to new subjects, and the creation of entirely novel combinations while preserving both the original subject and style.
@@
The recent surge of diffusion models [13, 36, 28] has unlocked previously unprecedented automatic image generation capabilities. The primary means of controlling such models is through text prompts, but text-based conditioning is inadequate for capturing the details of nuanced concepts, such as specific object instances (my dog) or individual styles (my child’s crayon drawing). Example-driven generation is highly desired in such scenarios. To this end, model fine-tuning methods like DreamBooth [30] and StyleDrop [34] capture subject or style from reference image(s) in a way that allows for novel renditions. However, such approaches tend to focus on either content or style in isolation and cannot be easily made to capture both or perform disentanglement. On the other hand, stylization techniques like ZipLoRA [33] and B-LoRA [6] can combine subject and style, but require the training of two separate models using subject- and style-specific input images. However, the scenario we consider requires the opposite – decomposing the subject and style from a single image.
@@
In this work, we introduce UnZipLoRA, a novel method that deconstructs an image into its constituent subject and style, represented as two distinct LoRAs (Low-Rank Adaptations [14]) trained simultaneously. These LoRAs can be used independently to generate variations of the subject or style and allow for recontextualizations. Moreover, our joint training method ensures that the resulting LoRAs are inherently compatible, i.e., can be seamlessly combined by direct addition to reconstruct the original image or to generate novel compositions of subject and style while preserving their fidelity. In fact our approach can be seen as a next stage of concept extraction – a problem previously studied in concept decomposition methods like Inspiration Tree [38] and CusConcept [41] that rely on textual inversion [8] to learn multiple text embeddings corresponding to the subconcepts within a set of images. However, textual inversion alone, without fine-tuning of weights, does not provide adequate control over or fidelity of the extracted concepts, which tend to remain generic and fail to capture nuances of the input object/style.
@@
Our key challenge is to learn two independent LoRAs simultaneously using only a single input image as supervision while ensuring that the resulting LoRAs correctly capture the subject and style concepts. Typical LoRA fine-tuning operates by binding the LoRA weights with the trigger phrase representing a specific subject or style concept in the input image – such as “a <cc> in <ss> style”. If we apply such a naive approach in our case, the presence of both the trigger phrases <cc> and <ss> in a single input prompt makes it difficult for the subject and style LoRAs to bind to the correct trigger phrase, resulting in cross-contamination. To solve this problem, we propose a novel prompt separation strategy that uses different prompts for each LoRA and the base model, and then combines them together in the intermediate feature space of the diffusion model in such a way that the loss for each LoRA can be calculated jointly using only the input image as supervision.
While prompt separation allows for joint training of LoRAs, the resulting LoRAs may not be compatible with each other, i.e., combining them through direct addition may produce poor quality recontextualizations. To make them inherently compatible, we also propose column separation and block separation strategies. In particular, column separation determines the importance of each column of LoRA weight matrix and adaptively assigns each column to either a subject LoRA or style LoRA using a dynamic importance re-calibration strategy. Such disjoint assignment ensures that high-importance columns from each LoRA remain decoupled. Block separation reserves some blocks of the U-net predominantly for style or for the subject, providing a further degree of disentanglement.
We demonstrate the effectiveness of UnZipLoRA for accurate separation of content and style using both human studies and quantitative metrics. Our results show a clear advantage of our method over separate LoRA fine-tuning via DreamBooth [30], concept separation via Inspiration Tree [38], and even the most recent state-of-the-art B-LoRA method [6]. We also showcase our method’s ability to preserve the concept fidelity for a wide array of recontextualizations – whether for using subject or style separately or together. In addition, our method provides the valuable capability for cross-composition of subject and style LoRAs obtained from different images. Finally, we demonstrate generalizability of our approach by providing results on KOALA – a newer, more efficient text-to-image model [19].
@@ 2 Related Work
Fine-tuning diffusion models is an effective way to personalize the text-to-image (T2I) models to depict specific concepts based on textual descriptions. Textual Inversion [8] optimizes the text embedding to represent a specific visual concept. DreamBooth [30] fine-tunes the diffusion model itself to better capture an input concept from a small number of images, while another group of methods [10, 17] aim to optimize specific parts of the networks to capture visual concepts. Most personalization approaches have quickly adopted LoRA [14], a fine-tuning technique that only optimizes a small subset of weights by low-rank approximations, as it is efficient for training and can mitigate overfitting problems.
@@
Image stylization is an area of research dating back at least 20 years [5, 12]. Great advances in arbitrary style transfer were achieved by approaches based on convolutional neural networks [9, 16, 15, 20, 24]. With the advent of deep generative models, a variety of approaches have attempted to fine-tune a pre-trained Generative Adversarial Network (GAN) or diffusion model for stylization [21, 7, 44, 23, 39, 18, 3, 32, 43, 42, 34, 4]. While these works provide valuable insights into style learning using generative models, the task we attempt to solve is the opposite: instead of stylizing a content image, we attempt to decompose a stylized image into its subject and style.
@@
Content-style decomposition. The task of subject-style decomposition can also be seen as a type of concept extraction. Inspiration Tree [38] aims to learn multiple embeddings corresponding to hierarchical subconcepts within a set of images. However, its reliance on textual inversion [8] limits this method to primarily producing text embeddings rather than the full LoRA weights needed for granular control over generation. Similarly, CusConcept [41] decomposes an image into visual concepts by learning customized embeddings. However, it does not produce dedicated LoRAs for each concept. Its reliance on computationally expensive LLMs and lack of explicit content-style separation limits its practical applicability. U-VAP [40] is a fine-tuning method that allows users to specify desired attributes from a set of images, enabling the disentangled use of visual concepts in diverse settings. However, this method requires elaborate data augmentation aided by LLMs, posing scalability challenges. ConceptExpress [11] explores unsupervised concept extraction and recreation by leveraging the inherent capabilities of pre-trained diffusion models. However, its reliance on localized masks extracted from the U-Net limits its application to concepts that can be localized within an image. This constraint prevents it from effectively handling global concepts like style. This limitation applies to Break-A-Scene [1] too, which relies on localized masks.
@@
The method most directly aimed at content and style separation is B-LoRA [6]. The authors of B-LoRA analyze the architecture of the SDXL base model [26] to find U-Net blocks most responsible for capturing content and style. By independently training content and style LoRAs restricted to the respective blocks, they obtain models that can be successfully mixed for various stylization applications. By contrast, UnZipLoRA trains both the LoRAs simultaneously. In our attempt to take advantage of B-LoRA’s block constraints for subject-style separation, we find that they are coarse-grained and thus insufficient to disentangle the subject in joint training. Therefore, we extend the block separation to more blocks and perform more fine-grained disentanglement. Together with our improved block separation strategy, the prompt separation strategy of UnZipLoRA provides a significant improvement in disentanglement ability, while column separation further enhances the fidelity of fine details.
@@ 3 Method
Figure 2: Method Overview. (1): Our training pipeline takes three prompts to learn two separate LoRAs. The column masks are introduced to establish the orthogonality. (2)-(4): Inference with individual learned LoRA or the combined LoRA.
@@ 3.1 Preliminaries
Diffusion models [13, 36, 28] are state-of-the-art deep generative models renowned for their ability to synthesize high-quality photorealistic images. In this work, we focus on Stable Diffusion XL (SDXL) [26], a widely used U-Net-based latent diffusion model (LDM) [28] known for its strong performance. In Section 4.5, we also show results on the newer, more efficient KOALA model [19].
Low-Rank Adaptation (LoRA) is an efficient fine-tuning method for adapting large vision or language models to new tasks [14, 31]. LoRA assumes that weight updates $\Delta W$ during fine-tuning have a low intrinsic rank and can thus be decomposed into two low-rank matrices, $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$, such that $\Delta W=BA$, where $r$ is the intrinsic rank of $\Delta W$ with $r<<\min(m,n)$. During training, only $A$ and $B$ are updated while keeping the original weights $W_{0}$ constant. The updated weights become $W=W_{0}+BA$. In case of text-to-image LDM models, model customization can be achieved by using LoRA fine-tuning to minimize reconstruction loss $\mathcal{L}_{DB}$ as proposed by DreamBooth [30].
@@ 3.2 Problem Setup
We aim to extract content and style from a single input image by learning two distinct LoRAs simultaneously. Given a pre-trained diffusion model with weights $\{W_{0}^{i}\}$, we learn two models: content LoRA $L_{c}=\{\Delta W_{c}^{i}\}$ and style LoRA $L_{s}=\{\Delta W_{s}^{i}\}$. Here, $i$ denotes the index of layers of the diffusion U-Net. In the following, we will skip the superscript $i$ as we operate over all LoRA-enabled weights of the base model. Once trained, the resulting LoRAs can be used either separately or together to achieve various recontextualizations as depicted in Fig. 2.
We follow the standard prompt construction strategy “a <cc> in <ss> style” with trigger phrases <cc> and <ss> to describe the content and style respectively [30, 34, 6, 17, 33]. Following prior works such as DreamBooth [30], the subject trigger phrase <cc> is formed with a unique token followed by the subject class label (e.g., ‘sks dog’), and the style trigger phrase <ss> consists of a generic description of style such as ‘watercolor painting’ [33, 34, 6]. We find that a single-word class label for the subject, and a high-level, brief (2-3 word) description for style are sufficient to effectively guide our method (see the supplementary material for further discussion of trigger phrase selection).
@@ 3.3 UnZipLoRA
In this section, we present our UnZipLoRA approach, which relies on three key components to ensure accurate disentanglement. First, prompt separation (Sec. 3.3.1) allows for training the subject and style LoRAs simultaneously. Second, column separation (Sec. 3.3.2) uses a dynamic importance recalibration strategy to make the columns of resulting LoRAs mutually orthogonal. Third, we introduce block separation (Sec. 3.3.3) that reserves certain style/subject-sensitive blocks of the SDXL U-Net only for style/subject learning to improve the fine details.
@@ 3.3.1 Prompt Separation
Our core challenge lies in the utilization of a single image as supervision for training two separate LoRAs. Since the input image contains both subject and style, we must use both LoRAs $L_{c}$ and $L_{s}$ together to calculate the loss during training. If we also use <cc> and <ss> in the input prompt, as in “A <cc> in <ss> style”, the training would lead to cross-contamination since the tokens corresponding to the style descriptor <ss> would be attended to by the cross-attention layers of the subject LoRA, and vice versa. This is evident in the Baseline (DreamBooth-LoRA) row of Fig. 5, where both the subject and style LoRAs are overfitted to input image.
@@
If we add the content and style LoRAs to the base model and use both <cc> and <ss> in the prompt $x$, the mapping in the cross-attention layer becomes
@@
where $x$ is the embedding for the combined prompt “A <cc> in <ss> style”, while $x_{c}$ and $x_{s}$ are the embeddings of subject and style descriptors <cc> and <ss> respectively. This allows each LoRA to attend to different concepts. As illustrated in Fig. 5 (row M1), adding prompt separation to a DreamBooth baseline prevents cross-contamination and successfully destylizes the content.